New Flyer Industries has decided to expand its production of hybrid transit buses. The firm expects incremental cash flows of $40 million per year for the next 10 years. The upfront cost of the expansion is $150 million and there are additional issuance costs for external financing of $15 million. If the New Flyers WACC is 7.5%, what is the NPV of the project? O A. S125 million O B. S110 million O C. $219 million O D. $95 million O E. $235 million
Answer #1

Net present value = Present value of cash outflows - Present value of cash outflow

Here the New flyer industries has incurred upfront costs of $ 150 million and additional cost of $ 15 million. Therefore, the total cash outflow is $ 150 + 15 million = $ 165 millions and there is no need to calculate present value as it is incurred presently.

Secondly, the present value of cash inflow = Annual cash inflow * PVAF(r%,t)

Where Annual cash inflow = $ 40 million per year
Rate = 7.5% and time is 10 years.

Therefore, Present value of cash inflow = 40 million * PVAF ( 7.5%, 10)
Present value of cash inflow = 40 * 6.86408 = $ 275 million (approximately)

Net present value = Present value of cash outflows - Present value of cash outflow

Net present value = $ 275 million - $165 million = $110 million

So the answer is b part i.e. $ 110 millions